OpenStudy (anonymous):

How many solutions does the following system have? { 16x - 8y=20 { 4x - 2y= 5 A. 0 B. 1 C. 2 D. Infinitely many

OpenStudy (anonymous):

(1) 16x - 8y = 20 (2) 4x - 2y = 5 Multiply equation 2 by -4: (1) 16x - 8y = 20 (2) -16x + 8y = -20 Add the equations: 0x + 0y = 0 0 = 0 Therefore, there are infinitely many solutions.